# Why do Learning Rates Transfer? Reconciling Optimization and Scaling Limits for Deep Learning

### v1

@@ 4 Case study: Two-Layer Linear Network
$|\lambda_{\max}[\gamma^{2}\nabla^{2}\mathcal{L}(E,V)]-\lambda_{\max}[\Theta(E,V)]|\leq\sqrt{\frac{\gamma}{ND}}\|w-w_{*}\|_{2}.$
The result above is actually more general: it indicates that the eigenvalues of the positive semidefinite portion of the Hessian $\mathcal{G}=K^{\top}K$ are fully [^7] Simple application of the SVD decomposition. characterized by the eigenvalues a smaller matrix $\Theta=KK^{\top}$. Furthermore, in our model $\Theta$ has a closed form that depends only on the variables $e,v$. In the next subsection, we characterize how $e,v$ evolve through time, and give conclusions for $\mu$P in Sec. 4.3.
4.1 Dynamics and Edge of Stability in Latent Space
In Sec. 4.3, we show at any value of the width $N$, under GD on the original network parameters $(E,V)$, the dynamics of $w,e,v$, can be described completely through a self-contained dynamical system in $(1+D+D^{2})$ dimensions. This property is surprising because the original dynamical system described by GD on the variables $E,V$ lives in $N(D+1)$ dimensions. Concretely, this means we can study the Hessian dynamics at different network widths in the same space. This result differs significantly in nature from the DMFT equations in (Bordelon and Pehlevan, 2022), where objects evolving are width dependent (evolving objects are features at different layers).
Theorem 4.2 (Evolution Laws).
Let $(E,V)$ evolve with GD at stepsize $\eta=\eta_{0}\gamma^{2}$ on the loss of Eq. 5. The evolution of $(w,e,v)$ is completely described by the following self-contained equation: let the ${}^{+}$ denote updated quantities,
While the system above describes the evolution laws $(w_{k},e_{k},v_{k})\to(w_{k+1},e_{k+1},v_{k+1})$, the dynamics are influenced also by initialization.
Proposition 4.3.
Last, by analyzing the stability of the dynamical system in Theorem 4.2, we can characterize the edge of stability using tools from dynamical systems (Ott, 2002) and Lemma 4.1.
@@ 4.1 Dynamics and Edge of Stability in Latent Space
Let $(E,V)$ evolve with GD with stepsize $\eta=\eta_{0}\gamma^{2}$ on the loss of Eq. 5 towards a minimizer ($E_{*},V_{*}$). Assume the corresponding solution in latent space $(w_{*},e_{*},v_{*})$ is marginally stable  [^8] In a dynamical system sense, i.e. some eigenvalues of the Jacobian have unit norm, while others have norm $<1$.. Then,
$\lambda_{\max}[\gamma^{2}\nabla^{2}\mathcal{L}(E_{*},V_{*})]=\frac{2}{\eta_{0}}\pm\frac{\eta_{0}\gamma^{2}\|w_{*}\|^{2}}{ND}.$
4.2 Implications for NTP
Consider $\gamma=1$ in Thm. 4.2. The dynamics of $(w,e,v)$ are width-dependent. Let us take $N\to\infty$ in the equation above to amplify this effect: the system becomes linear
@@ 4.2 Implications for NTP
While $w$ evolves from $w_{0}$ as expected from standard NTK theory (Jacot et al., 2018), $e,v$ stay clamped at initialization. Together with Lemma 4.1, this proves sharpness has no dependency on the learning rate in the width limit (we observe this e.g. in Fig. 1 and throughout all our experiments). This derivation is also in line with our discussion in Sec. 3.1: we only have sharpening under feature learning, and for the same reason we cannot observe NTP at the edge of stability as $N\to\infty$ (see stepsize dependency in Thm. 4.4), as also noted empirically by (Cohen et al., 2021).
4.3 Implications for $\boldsymbol{\mu}$P
The following result immediately follows by inspection of the equations in Thm 4.2, combined with Prop. 4.3.
@@ 4.3 Implications for $\boldsymbol{\mu}$P
Consider $\mu P$ ($\gamma=\sqrt{N}$) and let $(E,V)$ evolve with GD with stepsize $\eta=\eta_{0}\gamma^{2}$ on the loss of Eq. 5. Then, the equations governing the evolution of $(w,e,v)$ (defined in Thm. 4.2) in latent space have no width dependency – this holds at any finite width and not just at the limit. Initialization of $(w,e,v)$ is instead width-dependent, yet the error from $N\to\infty$ case scales in expectation like $1/\sqrt{N}$.
The corollary shows that $\mu$P trajectories at different widths align in the latent space $(w,e,v)$, albeit [^9] At moderate value of $N$, an initialization seed variation is at the same fluctuation scale as a change in width (see Prop 4.3). a vanishing perturbation in the initial condition. This feature of $\mu$P, especially when compared to the NTP setting in Sec. 4.2: while NTP’s dynamics for $e$ and $v$ become slower as the width increases, for $\mu$P their evolution laws are width independent (Cor. 4.5). This becomes especially important at the edge of stability: Lemma 4.1 shows that, as we approach a solution at the edge of stability, the sharpness for the original loss is largely determined by the value of $e$ and $v$ since (1) by Lemma 4.4 being at the edge of stability requires a sizeable sharpness value on the order $O(2/\eta_{0})+O(\eta_{0})$ and (2) the residual $\|w-w^{*}\|$, bounding the approximation error between $\lambda_{\max}[e+v\cdot I_{D}]$ and the true sharpness, becomes small (Fig. 5). These findings are in line with the general picture presented in Sec. 3.1.
@@ 5 Discussion & Conclusions
In this paper, we have provided insights into the sharpness evolution in optimization for neural networks by comparing the sharpness dynamics under different scaling limits. We have also argued how our results can explain the phenomenon of hyperparameter transfer under particular scaling limits, in that the sharpness is largely independent of the width under the scaling that achieves successful transfer. Finally, we have attributed the cause of the different sharpness evolution under different parametrizations to feature learning through the dynamics of the NTK matrix, thus drawing a connection between optimization and feature learning in deep neural networks.
Our results on the early training dynamics complement the analysis of Vyas et al. (2023) and Kalra et al. (2023) on the consistency of the loss and the sharpness in the first few steps of training. However, we further extend it, showing how while the loss curves depart later in training (with “wider/deeper being better"), the consistency of the sharpness’ dynamics is maintained longer in time. Furthermore, our work explains some of the experimental results on the lack of progressive sharpening under the NTP parameterization Cohen et al. (2021) (Appendix H), by relating it to the lack of feature learning and the consequent absence of hyperparameter transfer. Our results are also in line with the recent experiments on the evolution of the sharpness for ReLU MLPs trained with gradient descent under $\mu$P Kalra et al. (2023) (e.g. Figure 1). We extend these results to include the width (in)-dependence behaviour of the convergent stability point, a crucial aspect for successful hyperparameter transfer.
Finally, it is worth mentioning that there exist a family of infinite depth limits that admit feature learning ($\alpha\in[1/2,1]$ in Eq. 1, with an appropriate depth correction to the learning rate). Most of our experiments focus on models with a single layer per residual block, which exhibit transfer more consistently under the $\sqrt{L}-\mu$P setting (i.e. $\alpha=1/2$) and it is considered optimal in terms of feature diversity across blocks Yang et al. (2023). We adopt the same depth scaling with our experiments with Transformers — which have multiple layers per block. Under this setting, both Bordelon et al. (2023) (e.g. Fig. 3) and Yang et al. (2023) (Fig. 16) show good transfer across depth with Adam. More broadly, we expect that a scaling limit that admits learning rate transfer would find a corresponding width/depth-independent behaviour in the sharpness dynamics.
Future directions. Our results show that a width-independent sharpness holds consistently for large part of training. A theoretical study of the convergence rate of the dynamics of Hessian-related quantities such as its trace or sharpness would shed more light on this phenomenon. We also leave it as future work to investigate the comparison between different feature learning limits Yang et al. (2023), to give a closer look into the necessity/sufficiency of feature learning for hyperparameter transfer. More generally, we foresee that our paper could sparkle further research interest at the intersection between the scaling limits of neural networks and optimization theory.
6 Acknowledgments
@@
The authors would like to thank Bobby He, Imanol Schlag and Dayal Kalra for providing insightful feedback on early versions of this manuscript. LN would also like to thank Blake Bordelon, Boris Hanin and Mufan Li for the stimulating discussions on the topic of scaling limits that helped inspiring this work. AO acknowledges the financial support of the Hector Foundation.
@@ Appendix A Related works
It is worth mentioning that while for width limits feature learning is a clear discriminant between NTK and mean-field limits, the issue becomes more subtle with depth limits. In fact, there exist a family of infinite depth limits that admit feature learning ($\alpha\in[1/2,1]$ in Eq. 1, with an appropriate depth correction to the learning rate). Yang et al. (2023) classifies the depth limits in terms of the feature diversity exponent, a measure that quantifies the diversity between the features of different layers. With respect to this measure, $\alpha=1/2$ is the one that maximizes it. Bordelon et al. (2023) try to quantify the finite size approximation to the infinite (continuous) model, arguing that hyperparameter transfer is achieved faster for models that have lower discretization error to the infinite model’s dynamics.
Appendix B Experimental details
The implementations of our models are done in PyTorch. For the sharpness measurements, we use PyHessian Yao et al. (2020). We measure the sharpness on the same fixed batch throughout training. We perform selected experiments with data augmentations, where the random transformations are compositions of crops, horizontal flips, and $10$-degree rotations. Additionally, we provide further details on the models used in our experiments and the modifications we have introduced.
B.1 GPT-2
Our implementation is based on the implementation provided by Yang et al. (2022), with the addition of the residual scaling. Similar to the experiments performed on ViTs, the GPT-2 models are trained using Adam, where the base width is fixed and the depth is varied. In addition, we place the layer normalization layer in front of the residual, following the Pre-LN architecture.
B.2 Vision Transformers (ViTs)
The ViT implementation is based on the work of (Dosovitskiy et al., 2020) and follows the same tokenization and training protocol. In order to follow the Depth-$\mu$P parameterization, we make the same modifications as in B.1. The models are trained with Adam.
B.3 ResNet
We use convolutional neural networks with skip connections, with $3\times 3$ kernels and stride $1$. We apply pooling after every $3$rd layer, followed by a subsequent convolutional layer. Following Bordelon et al. (2023), Yang et al. (2023), we downscale the residual branches by $1/\sqrt{L}$.

### v2

@@ 5.1 Dynamics and Edge of Stability in Latent Space
$|\lambda_{\max}[\gamma^{2}\nabla^{2}\mathcal{L}(E,V)]-\lambda_{\max}[\Theta(E,V)]|\leq\sqrt{\frac{\gamma^{2}}{ND}}\|w-w_{*}\|_{2}.$
We stress that this result implies that evolution of the NTK (i.e. feature learning) goes hand in hand with the evolution of the sharpness, as we empirically show in Sec. 4.1. We are now ready to state the result on the sharpness at convergence.
@@
Let $(E,V)$ evolve with GD with stepsize $\eta=\eta_{0}\gamma^{2}$ on the loss of Eq. 4 towards a minimizer ($E_{*},V_{*}$). Assume the corresponding solution in latent space $(w_{*},e_{*},v_{*})$ is marginally stable  [^3] In a dynamical system sense: some eigenvalues of the Jacobian have unit norm, others have norm $<1$.. Then, $\lambda_{\max}[\gamma^{2}\nabla^{2}\mathcal{L}(E_{*},V_{*})]\in\left[\frac{2}{\eta_{0}},\frac{2}{\eta_{0}}+\frac{\eta_{0}\gamma^{2}\|w_{*}\|^{2}}{ND}\right].$
Implications for NTP.
Consider $\gamma=1$ in Thm. 5.1. The dynamics of $(w,e,v)$ are width-dependent. Let us take $N\to\infty$ in the equation above to amplify this effect: the system becomes linear
@@
While $w$ evolves from $w_{0}$ as expected from standard NTK theory (Jacot et al., 2018), $e,v$ stay clamped at initialization. Applying Lemma 5.2 with $\gamma=\mathcal{O}(1)$, the Hessian and the NTK have the same largest eigenvalue at large width (at rate $O(\sqrt{N})$). This agrees with our intuition, as under NTP the predictor converges to a linear model in the large $N$ limit, and thus $\mathcal{R}$ vanishes. Also, this proves that the sharpness has no dependency on the learning rate in the width limit (we observe this, e.g., in Fig. 1 and throughout all our experiments). This derivation is also in line with our discussion in Sec. 4.1: we only have sharpening under feature learning, and for the same reason we cannot observe NTP at the edge of stability as $N\to\infty$ (see stepsize dependency in Prop. 5.3), as also noted empirically by (Cohen et al., 2021).
Implications for $\boldsymbol{\mu}$P.
The following result immediately follows by inspection of the equations in Thm 5.1, combined with Prop. C.1.
@@
Consider $\mu P$ ($\gamma=\sqrt{N}$) and let $(E,V)$ evolve with GD with stepsize $\eta=\eta_{0}\gamma^{2}$ on the loss of Eq. 4. Then, the equations governing the evolution of $(w,e,v)$ (defined in Thm. 5.1) in latent space have no width dependency – this holds at any finite width and not just at the limit. Initialization of $(w,e,v)$ is instead width-dependent, yet the error from $N\to\infty$ case scales in expectation like $1/\sqrt{N}$.
The corollary shows that $\mu$P trajectories at different widths align in the latent space $(w,e,v)$, albeit with a vanishing perturbation in the initial condition (see Prop. C.1). While NTP’s dynamics for $e$ and $v$ become slower as the width increases, for $\mu$P their evolution laws are width independent. This implies that if the dynamics converge towards a minimizer for $e$ and $v$, this will be at the sharpness value predicted by Prop. 5.3. Under $\mu$P, where $\gamma^{2}\propto N$, this value will be width-independent, as Super Consistency would suggest. We stress that Prop. 5.3 characterizes the sharpness at convergence (i.e. at infinite time). At finite time, there is still a discrepancy between the $\lambda_{\max}(\Theta)$ and the sharpness of the order of the residual term $1/\sqrt{D}\|w-w^{*}\|$ (Lemma 5.2). Finally, we stress that Prop. 5.3 prescribes the right scaling for the Hessian by including the preconditioning factor of $\gamma^{2}$. Thus, we do not prove that at any finite time, the whole sharpness trajectory is width-independent, nor we are estimating converge rates in $N$ at finite time. Indeed, there will be a finite-size dependence coming from the initial conditions. We leave a precise characterization of the whole sharpness dynamics across the training trajectory for future work.
@@ 6 Discussion & Conclusions
On Feature Learning Parametrizations. In this paper, we have shown how certain properties of the loss Hessian evolve almost identically across training for different model sizes, and named this property Super Consistency. We have also compared the sharpness dynamics under different scaling limits and parameterizations, and related Super Consistency of the landscape to learning rate transfer. Beyond being able to distinguish feature learning (rich) and kernel (lazy) parametrization, we have also shown how other suboptimal feature learning parametrizations have sharpness dynamics violating Super Consistency through finite-size accumulations. This seems to suggest that Super Consistency of the landscape is an important discriminant when it comes to hyperparameter transfer beyond the rich/lazy regimes. We foresee that our paper could spark further research interest at the intersection between the scaling limits of neural networks and optimization theory.
On the NTK and Hessian dynamics. In Section 4.1 we have drawn the connection between progressive sharpening and NTK evolution in the early phase of training. However, Figure 3 (b), shows how the NTK eigenvalues at different widths accumulate finite-size effects over time and diverge from each other, while the Hessian eigenvalues are Super Consistent. This suggests that other forces are at play after progressive sharpening, such as Self-Stabilization Damian et al. (2022). In fact, progressive sharpening on one hand, and Self-Stabilization on the other, make the stability threshold a stable attractor of the sharpness dynamics. Gaining theoretical understanding for these complex interactions in the context of scaling limits is an exciting area of future research.
Design of Step-size Tuners. In most of our experiments, we rely on a constant step size $\eta_{0}$. However, an alternative is to use a step-size tuner, i.e. to automatically choose $\eta_{0}$ based on some criteria of the local landscape (Roulet et al., 2023). Our results open directions into some possible investigations and design choices for new step size tuners. For instance, do step size tuners transfer with the width and depth of the architecture? Given our results on the role of warmup schedule to improve transfer, it seems plausible to design step size tuners that use EoS results to achieve optimal learning rate transfer under different parameterizations.
Limitations. One of the underlying assumptions of the argument presented here is that the sharpness is an important property of the landscape when it comes to step size selection. Indeed, the results in Cohen et al. (2021) establish a more intricate relationship between sharpness and learning rate. We discuss this in Sec. A.2. Overall, our theory on Super Consistency does not exclude the existence of other factors that might influence the optimal learning rate. Hyperparameter transfer requires Super Consistency of the landscape, thus we expect other potential factors to have this property. Finally, we note that due to the high computational cost of Hessian estimation, we do not perform experiments at a larger scale than presented here. It would be interesting to see if Super Consistency still holds at an even larger scale.
Acknowledgements
@@
The authors would like to thank Bobby He, Imanol Schlag, Dayal Kalra, Tiago Pimentel and Gregor Bachmann for providing insightful feedback on early versions of this manuscript. LN would also like to thank Blake Bordelon, Boris Hanin and Mufan Li for the stimulating discussions on the topic of scaling limits that helped inspiring this work. AO acknowledges the financial support of the Hector Foundation. LN acknowledges the support of a Google PhD fellowship. AM acknowledges the support of a Kempner Fellowship.
Appendix
[…]
@@ Appendix A Related works
It is worth mentioning that while for width limits feature learning is a clear discriminant between NTK and mean-field limits, the issue becomes more subtle with depth limits. In fact, there exist a family of infinite depth limits that admit feature learning ($\alpha\in[1/2,1]$ in Eq. 1, with an appropriate depth correction to the learning rate). Yang et al. (2023) classifies the depth limits in terms of the feature diversity exponent, a measure that quantifies the diversity between the features of different layers. With respect to this measure, $\alpha=1/2$ is the one that maximizes it. Bordelon et al. (2023) try to quantify the finite-size approximation to the infinite (continuous) model, arguing that hyperparameter transfer is achieved faster for models that have lower discretization error to the infinite model’s dynamics.
A.1 Learning Rate Transfer and Scaling Limits
Our results on the early training dynamics complement the analysis of Vyas et al. (2023) and Kalra et al. (2023) on the consistency of the loss and the sharpness in the first few steps of training. However, we further extend it, showing how while the loss curves depart later in training (with “wider/deeper being better"), the consistency of the sharpness’ dynamics is maintained longer in time and does not accumulate finite-size effects over time. Furthermore, our work explains some of the experimental results on the lack of progressive sharpening under the NTP parameterization Cohen et al. (2021) (Appendix H), by relating it to the lack of feature learning and the consequent absence of hyperparameter transfer. Our results on the role of feature learning are also compatible with Agarwala et al. (2022), where it is theoretically shown that the non linear dynamics exhibits progressive sharpening in a simple non linear model. Our results are also in line with the recent experiments on the evolution of the sharpness for ReLU MLPs trained with gradient descent under $\mu$P Kalra et al. (2023) (e.g. Figure 1). We extend these results to include the width (in)-dependence behaviour of the convergent stability point, a crucial aspect for successful hyperparameter transfer.
Finally, It is worth mentioning that there exist a family of infinite depth limits that admit feature learning ($\alpha\in[1/2,1]$ in Eq. 1, with an appropriate depth correction to the learning rate). Most of our experiments focus on models with a single layer per residual block, which exhibit transfer more consistently under the $\mu$P- $\sqrt{\text{depth}}$ setting (i.e. $\alpha=1/2$) and it is considered optimal in terms of feature diversity across blocks Yang et al. (2023). We adopt the same depth scaling with our experiments with Transformers — which have multiple layers per block. Under this setting, both Bordelon et al. (2023) (e.g. Fig. 3) and Yang et al. (2023) (Fig. 16) show good (albeit at time slightly worse) transfer across depth with Adam. More broadly, we expect that a scaling limit that admits learning rate transfer would find a corresponding width/depth-independent behaviour in the sharpness dynamics.
A.2 Sharpness and Optimal Step Size
One of the underlying assumptions to explain why super consistent sharpness causes hyperparameter transfer is that the sharpness influences the optimal learning rate. Although this is well-established in classic optimization LeCun et al. (2002), Edge of Stability tells a story of a more intricate relationship: the choice of the learning rate also influences the sharpness. Also recent works on step size tuners show evidence of a more complex interaction in the joint dynamics of step size and sharpness Roulet et al. (2023), a relation that still has to be fully understood and could be leveraged to design better step size tuners. However, the sharpness is still arguably a very good proxy for understanding the trainability of the model at large learning rates. For instance, Gilmer et al. (2021) argue that maintaining a small sharpness in neural network optimization favors training at large learning rates. Interestingly, in Cohen et al. (2021) (Appendix F) it is shown that choosing the step size as $1/\lambda_{t}$ at any step is suboptimal. The reason could be that gradient directions are not aligned with the largest curvature. Alternatively, one could claim there is another, more sophisticated, functional relationship between sharpness and the maximum step size allowed. Indeed, the “catapult mechanism”Lewkowycz et al. (2020) shows the maximum stable learning rate is $c_{arch}/\lambda_{0}$, where $c_{arch}$ depends on the architecture. On a related note, Gur-Ari et al. (2018) shows that after the first few steps, the gradient direction is aligned with the top Hessian eigenvectors, underlying once again the importance of the sharpness and the first few eigenvalues in this phase of training, which we show here to be super consistent (Fig. 2).
Appendix B Experiments in the absence of a valid scaling limit
B.1 $\mu$P without $1/\sqrt{\text{depth}}$ scaling of the residual branches
We first study the effect of increasing the depth in a model without the $1/\sqrt{L}$-scaling of the residual branches, which results in exploding activations (and sharpness) as the depth increases. See Figure 6, where we first show that (top row) there is no learning rate transfer, and there is no consistent sharpness either. For instance, notice how the sharpness for the model of depth 24 quickly reaches its EoS value at its optimal learning rate of 0.068 (and larger models, e.g. depth 36 diverge). On the other hand, for the same learning rate the smaller depth models struggle to reach EoS at the same speed. This suggests that for the smaller-depth model, the learning rate can be safely increased. Indeed, the optimal learning rate for the smaller depth model is significantly larger, where the model reaches the EoS value very fast (optimal lr: 0.53). The larger depth models are not trainable at these larger learning rates.
In the bottom row of Figure 6, we show that by adding a linear warmup scheduler in the first phase of training, the network is progressively more trainable at larger step-sizes, and the sharpness reaches its stability threshold at any width/depth that is trainable Gilmer et al. (2021). This suggests that the diverging initial sharpness can be counteracted with initial small learning rates. Also, we notice a better transfer, indicating that a sustained period of depth-independent sharpness can help learning rate transfer. On the other hand we do not observe good transfer after the first epoch, which correlates with the fact that the sharpness is not consistent due to a blow up with depth at initialization.
Figure 6: $\mu$P without $1/\sqrt{\text{depth}}$ scaling of the residual branches for ConvNets trained using SGD on CIFAR10. (Top row): no warmup. (Bottom row): 200 steps of linear warmup. Right column: sharpness dynamics in the first epoch for the optimal learning rates for the models of depth 3 and depth 24. Notice how in the first epoch (Left column) there is no learning rate transfer in both cases. This correlates with the fact that that the training starts at an increasing sharpness with depth, which causes no consistency the dynamics (and divergent behaviour at large learning rates). Adding warmup alleviates this issue, allowing the model to reach edge of stability and improve learning rate transfer (middle column). One step in the sharpness plot corresponds to $2$ batches. Parameters: batch size$=128$, epochs$=20$, using data augmentation.
B.2 Standard Parameterization (SP) experiments
Following the SP formulation introduced in (Yang and Hu, 2021), we analyze the sharpness and learning rate transfer under this regime in Figure 7. Note that our experiments use a fixed learning rate $\eta$. While Yang and Hu (2021) use a width scaled learning rate in order to parameterize SP as a kernel limit in infinite width, the SP definition that we use does not have a well defined limit and thus diverges as the width is increased. For more details, we refer the reader to (Yang and Hu (2021), Appendix J.3).
Figure 7: Standard Parameterization (SP) for ConvNets trained using SGD on CIFAR10, for varying number of learning rate warmup steps. (Left column) No warmup, (Middle column) 1000 warmup steps and (Right column) 2000 warmup steps. Note that under SP, in the beginning the training starts from a high curvature, which means that large step sizes would lead to divergent behaviour. Adding warmup alleviates this issue, allowing the model to reach edge of stability and improve learning rate transfer. One step corresponds to $10$ batches. Parameters: batch size=$256$, epochs=$20$, using data augmentation.
B.3 $\mu$P - disabling residuals
We also provide training runs of models parameterized with $\mu$P, while disabling the residuals. Note that the models quickly become untrainable under this regime when increasing the depth. This phenomenon is due to the vanishing curvature experienced by these models, as shown in Figure 8, which leads to vanishing signal.
Figure 8: $\mu$P parameterization on ConvNets with residuals disabled ($\tau=0$) trained on CIFAR10. (Left) Learning rate transfer plot, showing that under this setting, the optimal learning rate transfers, but with increasingly larger shifts when increasing the depth due to the vanishing signal. (Right) Sharpness evolution during training, showing that the dynamics are following a depth independent trajectory for the optimal learning rate $(0.5275)$. Note that deeper models become much harder to train when residuals are disabled, motivated by the observation that the the depth $24$ model suffers from vanishing curvature at larger learning rates. The spikes in the plot are due to the curvature approaching 00 in log-scale. One step corresponds to $10$ batches. Parameters: batch size=$256$, epochs=$20$, using data augmentation.
B.4 $\mu$P experiments for full batch GD
In this section, we investigate the effect of $\mu$P and on learning rate transfer and sharpness, when trained with full batch gradient descent. We subsampled $5000$ sampled from CIFAR10 in a stratified fashion (i.e. $500$ sampled from each of the $10$ classes) and proceeded to train residual ConvNets under the $\mu$P parameterization with GD, following a similar procedure as (Cohen et al., 2021). Our findings show that under $\mu$P, when using a large enough learning rate, the models are able to achieve edge of stability, as well as have learning rate transfer. This is empirically shown in Figure 9, where we can see that while the sharpness has the typical oscillations around the EoS threshold studied by (Cohen et al., 2021), it still does maintain a width-independent trend during training.
Figure 9: Residual ConvNets trained using (full batch) GD on a $5000$ sample subset of CIFAR10. (Left) Learning rate transfer plot, showing that the optimal learning rate transfers across different widths; the slight shift in the transfer plot is due to the oscillations around the EoS threshold seen in (middle) and (right). (Middle) Sharpness dynamics during training, showing a consistent width independent dynamic throughout the whole training procedure; dashed line represents $2/\eta$. (Right) Training loss dynamic during time for the optimal learning rate $(0.76)$, showing the wider-is-better behaviour of the $\mu$P parameterization, as well as the oscillations induced by the EoS regime. Parameters: batch size$=256$, no warmup, using data augmentation.